NvidiaIntely (nvidiaintely):

You can study reaction rates to answer which of the following questions? Will the reaction go to completion? Will the reaction occur? How can the reaction be slowed down? How many grams of product will be made?

OpenStudy (kevin):

Maybe the answer is How can the reaction be slowed down

OpenStudy (kevin):

The formula for reaction time is : \[\frac{ 1 }{ t2 } = n^\frac{ \Delta T }{ a } \times \frac{ 1 }{ t1 }\] delta T is temperature change So we can slowed down the reaction by change the temperature value

NvidiaIntely (nvidiaintely):

Thanks kevin!

OpenStudy (kevin):

u are welcome :)